# Fractal geometry predicts dynamic differences in structural and functional connectomes

### v1

@@ Abstract
Understanding the intricate architecture of brain networks and its connection to brain function is essential for deciphering the underlying principles of cognition and disease. While, traditional graph-theoretical measures have been widely used to characterize these networks, they often fail to fully capture the emergent properties of large-scale neural dynamics. Here, we introduce an alternative approach to quantify brain networks that is rooted in complex dynamics, fractal geometry, and asymptotic analysis. We apply these concepts to brain connectomes and demonstrate how quadratic iterations and geometric properties of Mandelbrot-like sets can provide novel insights into structural and functional network dynamics. Our findings reveal fundamental distinctions between structural (positive) and functional (signed) connectomes, such as the shift of cusp orientation and the variability in equi-M set geometry. Notably, structural connectomes exhibit more robust, predictable features, while functional connectomes show increased variability for non-trivial tasks. We further demonstrate that traditional graph-theoretical measures, when applied separately to the positive and negative sub-networks of functional connectomes, fail to fully capture their dynamic complexity. Instead, size and shape-based invariants of the equi-M set effectively differentiate between rest and emotional task states, which highlights their potential as superior markers of emergent network dynamics. These results suggest that incorporating fractal-based methods into network neuroscience provides a powerful tool for understanding how information flows in natural systems beyond static connectivity measures, while maintaining their simplicity.
@@
Keyword1, Keyword2, Keyword3, Keyword4
@@ 1 Introduction
Fractal research transformed how we understand complexity, revealing how simple rules can generate intricate, emergent behavior. The Mandelbrot set – created by iterating a basic quadratic function – became an iconic example, illustrating self-similarity and nonlinear feedback in complex systems. Fractal geometry has long been recognized as a powerful framework for describing self-similar structures in nature, from vascular branching [1, 2] to neuronal morphology [3, 4].
Building on this legacy, we previously developed a generalizable framework for analyzing dynamics in complex networks. At its core is the equi-M set, a new object that extends classical complex dynamics into high-dimensional, evolving systems. We explore the role of equi-M sets as a tool for characterizing the global organization of brain networks. By mapping the asymptotic behavior of iterated quadratic dynamics onto empirical connectomes, we generate novel descriptors that integrate connectivity structure with properties of the emerging dynamics. Unlike traditional network analysis, which emphasizes static features such as node degree or clustering, our approach captures how behavior emerges over time through interactions across the network.
@@ 1.1 Complex quadratic dynamics
For iterations of a single quadratic map, one of the most striking results is that the trajectory (or orbit) of the critical point $z=0$ under iterations of the map captures global information on the behavior of all other orbits. The property of the critical orbit being bounded is particularly important in the context of long-term iterations. The Mandelbot set appeared in the 1960s from the need to obtain an overarching description of this property over the whole family of quadratic maps. It is defined as the set of all $c\in\mathbb{C}$ for which the critical orbit of the complex quadratic function $f_{c}(z)=z^{2}+c$ is bounded. This is a fractal with fascinating characteristics that have been extensively analyzed since its introduction. The topological organization of the Mandelbrot set into hyperbolic bulbs (leading to its shape and fractal geometry) classifies all possible combinatoric behaviors that orbits can have in the family $f_{c}$. This makes the Mandelbrot set both a staple in discrete dynamical systems and a canonical representation for more general phenomena.
@@ 1.2 Complex quadratic maps on networks
Our work extends this framework from its original scope of single map iterations to the broader context of complex dynamical networks (CGNs). More specifically, we consider networks in which the coupled nodes are identical complex quadratic maps, with coupling specified by the binary adjacency matrix $A\in\mathbb{R}^{n\times n}$ and weights $g\in\mathbb{R}^{n\times n}$. The overall evolution of this discrete system takes the form of an iteration in $\mathbb{C}^{n}$, which acts in each component as
@@
Since nodes have the same intrinsic dynamics dictated by the parameter $c$, the emerging network dynamics are a direct reflection of the value of $c$ and of the specific $(A,g)$ network architecture at play. By analogy to the single map case, we define the equi-M set of the network as the set of all $c\in\mathbb{C}$ for which the critical point $0\in\mathbb{C}^{n}$ remains bounded under iteration. In this generalized context, many of the results on the properties and structure of the traditional Mandelbrot set are lost. Our theoretical work shows that the equi-M set no longer provides a full combinatorial description of the asymptotic behavior of the map [12]. However, our results strongly suggest that equi-M sets have quantifiable topological and fractal properties that reflect the network architecture, making them good candidates for testing and classifying how dynamic patterns evolve for different architectural profiles. In other words, the equi-M set still captures enough of the long-term dynamic behavior to be used as an assessment of the network’s dynamic range.
@@ 2.1 Dataset description
Our analysis is based on tractography-derived and functional neural connectivity data obtained from the S1200 Q4 release of the Human Connectome Project (HCP) [13]. The public domain data contains extensive MRI, behavioral, and demographic scans from a large sample of individuals (> 1000). For this study, we considered a subgroup consisting of 48 arbitrarily chosen subjects.
@@ 2.2 Network measures
($BC$) is a global network measure of how often a node lies on the shortest path between all pairs of nodes in the network.
@@
($AS$) measures how likely nodes in a network are to connect with other nodes that have the same degree.
@@
($GE$) is the average (over all node pairs in the network) of the shortest path length between those two nodes.
@@
($TR$) is the ratio of “triangles to triplets” in the network [20].
@@
($M$) are patterns of network connections among a specified number of nodes forming a sub-graph $g$, which act as building blocks for complex networks [20]. Given that our Structural connectivity matrices are undirected, our analysis focused on two types of three-node motifs.
@@
Graph-theoretical measures for Functional connectomes
A common approach to analyzing signed networks, particularly in the context of brain connectomes, is to separate the positive and negative sub-networks and examine them independently. This method assumes that nodes exhibiting strong positive correlations (highly co-activated regions) and those with negative correlations (anti-correlated activity) form distinct functional communities. By analyzing these sub-networks separately, investigators aim to capture the intrinsic dynamics of cooperative and opposing activity patterns within the network [22]. However, this approach implicitly assumes that positive and negative sub-networks can independently encode key aspects of the network’s emergent behavior, an approach that has faced scrutiny [23, 24]. In this study, we aim to examine the validity of this assumption by exploring the extent to which overall network dynamics are influenced not only by these segregated sub-networks, but also by the intricate interplay between them. Therefore, for Functional networks, we separately compute the measures for the positive / negative sub-networks and also for the absolute value of the full functional connectome. We will call these the networks associated with the connectome.
Our findings show significant differences in the equi-M sets generated by the segregated sub-networks. This hints towards the fact that the interaction between the positively and negatively correlated regions contribute fundamentally to the emergent properties of the network, rather than simply being an artifact of its structural composition.
@@ 2.3 Construction of equi-M sets
The equi-M sets are computed numerically for the connectome of each of the 48 subjects using the netbrot visualizer [25]. Each group of equi-M sets is rendered in an appropriately scaled bounding box with a resolution of $1200\times 1200$ pixels. The iteration proceeds for a maximum of $512$ steps using an escape radius $R_{e}=100$ until it decides if the chosen point is in the set. The chosen parameters give reasonably accurate numerical representations for the theoretical equi-M set. Illustrations of the corresponding equi-M sets for all 48 subjects are included in Figures 12, 13 and 15, for all Structural and Functional connectomes that were used in this analysis.
@@
For each of the three types of connectomes (Structural, Rest, and Emotion) we computed statistical equi-M sets over all subjects in the data set. For each parameter point $c$ in the corresponding bounding box, we computed the fraction $f$ of the number of connectomes in the group, such that $c$ is in the equi-M set for that connectome. We illustrate these “stochastic” connectomes using pcolor plots, in which the color gradient (from white to black) represents the fraction $f$. We also took a mean-field approach to the data, and we generated equi-M sets for prototypical connectomes, computed as the mean of all connectomes in a data set (group average). These were defined in previous work [12] and are described in more detail in Section 3.
@@ 2.4 Geometric measures of equi-M sets
The equi-M sets for all Structural and Functional connectomes in our dataset are shown in Figures 12, 13 and 15 in order to allow a general view of their diversity in size and geometry and to facilitate direct visual comparisons. We use a variety of approaches to capture and quantify these visual differences and to understand their relationship to the network architecture.
@@
Figure 1: Measurable differences in equi-M set shape can be captured by computing a set of geometric measures, illustrated here for two random Structural connectomes in our data set (subjects #22 and #5, respectively). In the left panel, we have the following landmarks and distances: $C=0.13$, $T=-0.9$, $L=0.22+0.16i$, $R=-0.06+0.43i$, $d_{TC}=1.03$, $d_{TR}=0.94$, $d_{TL}=1.14$, $d_{CL}=0.19$, $d_{R\bar{R}}=0.41$, $\epsilon=0.87$. In the right panel: $C=0.08$, $T=-0.39$, $L=0.13+0.06i$, $R=-0.03+0.2i$, $d_{TC}=0.47$, $d_{TR}=0.42$, $d_{TL}=0.53$, $d_{CL}=0.008$, $d_{R\bar{R}}=0.87$, $\epsilon=0.84$.
@@
Figure 2: Measurable differences in equi-M set shape captured by computing geometric measures for sets derived from Functional connectomes. The two panels illustrate the different geometry landmarks and corresponding measures between two Emotion connectomes (subjects #28 and #3, respectively). In the left panel, we have the following landmarks and distances: $C=-0.01$, $T=0.024$, $S=0.025+0.001i$, $L=-0.021+0.003i$, $R=0.003+0.026i$, $d_{TC}=0.04$, $d_{SR}=0.03$, $d_{SL}=0.04$, $d_{CL}=0.006$, $d_{R\bar{R}}=0.05$, $\epsilon=1.3$. In the right panel: $C=-0.008$, $T=0.015$, $S=0.015$, $L=-0.013+0.007i$, $R=-0.0006+0.013i$, $d_{TC}=0.024$, $d_{SR}=0.021$, $d_{SL}=0.03$, $d_{CL}=0.008$, $d_{R\bar{R}}=0.02$, $\epsilon=1.1$.
@@
For each individual equi-M set, we computed a collection of measures that capture details of the 2-dimensional position, size, and shape of the set. To begin with, we noticed that the Structural sets (corresponding to symmetric connectomes with all positive entries) replicate some broad topological features of the traditional equi-M set. This is not surprising, and is explained in more detail in [12]. In particular, the sets are symmetric about the real axis and exhibit a main cardioid that delimit the locus of $c$ for which the critical point converges to an attracting fixed point. A feature that appears to be robust among all Structural connectomes is that the main cardioid has a cusp on the right, with varying position and depth. The main cardioid extends to the left with a “tail” formation. However, the tail structure is not robust and presents high variability of shapes. The one consistent feature remains that the tip of the tail is the leftmost point of the set. Therefore, we used the position $C$ of the cusp (with $\operatorname{Im}(C)=0$) and the position $T$ of the tip of the tail (with $\operatorname{Im}(T)=0$) as landmarks for the stretch of the set along the real axis. In addition, we considered the coordinates of the point $R$ farthest from the real axis (with $\operatorname{Im}(R)>0$) and the coordinates of the rightmost points (the lip of the cusp) $L$ (with $\operatorname{Im}(L)>0$). To encode the general shape, we obtained the horizontal diameter $d_{TC}=|T-C|$, the vertical diameter $d_{R\bar{R}}=2|\operatorname{Re}(R)|$, the depth of the cusp as $d_{CL}=|C-L|$, the slant diameters from the tail to the highest point $d_{TR}=|T-R|$ and to the lip of the cusp $d_{TL}=|T-L|$, and the “elongation” of the set, as the ratio $\epsilon=d_{R\bar{R}}/d_{TC}$. The collection of these measures is expected to encompass key information about variability in shape among Structural equi-M sets (as suggested by Figure 1, in which the measures are illustrated for two example sets.
@@
To identify similar landmarks for the Rest connectomes, we had to take into consideration the impact of the signed Functional connections on the topological structure. The first striking difference was that the orientation of the sets was “reversed” from that of the Structural connectomes. While all Rest sets still had a main cardioid body and a visually identifiable cusp, the cusp was left facing, and the (less prominent) tail formation was extending to the right. As before, we identified the cusp $C$, the tip of the tail $T$, the highest point $R$ and the upper lip of the cusp $L$. In addition, we identify a separate marker $S$, which is the rightmost point in the set, similarly to $L$ in the Structural equi-M sets. Together with this new marker, we define a slant diameter $d_{SL}=|S-L|)$ and the depth of the tail $d_{ST}=|S-T|$. We used this expanded set of measures to characterize the shape of the Rest sets.
@@
The Emotion sets presented with even more variability in the shape of the main body (see the complete illustration of all sets in Figure 15). Although most cardioids still had a left cusp, this was replaced in some cases by a more complex formation. Similarly, while most sets had a right tail, this was “stumpy,” and in some cases missing or replaced by another cusp. While in our theoretical work we are trying to analytically understand the conditions behind creating main cardioid cusps, this is beyond the scope of this paper. In this project, we simply use the same set of landmarks $(C,T,R,L,S)$ as for the Rest connectomes, and the same combinations between them, to broadly estimate the shape of the sets. However, in our Results and Discussion, we further comment on the significance of the differences in cusp and tail orientation and structure between the three types of connectomes.
@@
Our simulations suggest dramatic between-subject differences in the detail and variability of the equi-M set boundary, beyond the broad geometric measures described in the previous paragraphs. To capture these subtler properties, we constructed a truncated Fourier representation of the equi-M set boundary. For each equi-M set, the boundary was first identified and approximated using the Ramer–Douglas–Peucker algorithm [26] (part of the OpenCV library) to a specified tolerance of $10^{-4}P$, where $P$ represents an approximation of the set perimeter. The resulting polygonal boundary approximation was then used to obtain a truncated Fourier representation (as seen in Figure 3). Note that the boundary of an equi-M set is fractal in nature and its Fourier representation does not always converge in the limit. Even so, investigating the resulting Fourier modes (i.e. coefficients in the discrete Fourier expansion) can provide valuable geometric information about the sets.
We used the Fourier representation (see [27, Chapter 6]) of the boundary of the equi-M set to compute several additional measures. In general, the measures are invariant under translations, rotations, and (sometimes) scaling. Among these are the centroid $G$, the area $A$, and the perimeter $P$. Other measures, such as (average) curvature, are not representative due to the presence of cusps. Normalized Fourier modes (also known as Fourier descriptors) can also offer further insight into the large scale structure of the equi-M sets [27, Chapter 11].
@@
Figure 3: Fourier mode reconstruction of the Structural connectome from Figure 1a (top) and Functional connectome from Figure 2b (bottom) with $N+1$ modes in $\{-N/2,\dots,N/2\}$.

### v2

@@ Abstract
Understanding the intricate architecture of brain networks and its connection to brain function is essential for deciphering the underlying principles of cognition and disease. While traditional graph-theoretical measures have been widely used to characterize these networks, they often fail to fully capture the emergent properties of large-scale neural dynamics. Here, we introduce an alternative approach to quantify brain networks that is rooted in complex dynamics, fractal geometry, and asymptotic analysis. We apply these concepts to brain connectomes and demonstrate how quadratic iterations and geometric properties of Mandelbrot-like sets can provide novel insights into structural and functional network dynamics. Our findings reveal fundamental distinctions between structural (positive) and functional (signed) connectomes, such as the shift of cusp orientation and the variability in equi-M set geometry. Notably, structural connectomes exhibit more robust, predictable features, while functional connectomes show increased variability for non-trivial tasks. We further demonstrate that traditional graph-theoretical measures, when applied separately to the positive and negative sub-networks of functional connectomes, fail to fully capture their dynamic complexity. Instead, size and shape-based invariants of the equi-M set effectively differentiate between rest and emotional task states, which highlights their potential as superior markers of emergent network dynamics. These results suggest that incorporating fractal-based methods into network neuroscience provides a powerful tool for understanding how information flows in natural systems beyond static connectivity measures, while maintaining their simplicity.
@@
asymptotic dynamics, complex quadratic networks, brain network, cognitive task
Lead paragraph. Fractal research transformed how we understand complexity, revealing how simple rules can generate intricate, emergent behavior. The Mandelbrot set – created by iterating a basic quadratic function – became an iconic example, illustrating self-similarity and nonlinear feedback in complex systems. Building on this legacy, we developed a generalizable framework for analyzing dynamics in complex networks. At its core is the equi-M set, a new object that extends classical complex dynamics into high-dimensional, evolving systems. Unlike traditional network analysis, which emphasizes static features such as node degree or clustering, our approach captures how behavior emerges over time through interactions across the network. This dynamic perspective opens new pathways for understanding complex systems in both natural and engineered contexts.
@@ 1 Introduction
Fractal research transformed how we understand complexity, revealing how simple rules can generate intricate, emergent behavior [1]. Fractal geometry has long been recognized as a powerful framework for describing self-similar structures in nature, from vascular branching [2, 3] to neuronal morphology [4, 5]. In particular, the Mandelbrot set – created by iterating a basic quadratic function – became an iconic example, illustrating self-similarity and nonlinear feedback in complex systems [6, 7].
Building on this legacy, we previously developed a generalizable framework for analyzing dynamics in complex networks [8, 9, 10, 11, 12]. At its core is the equi-M set, a new object that extends classical complex dynamics into high-dimensional, evolving systems. We explore the role of equi-M sets as a tool for characterizing the global organization of brain networks. By mapping the asymptotic behavior of iterated quadratic dynamics onto empirical connectomes, we generate novel descriptors that integrate connectivity structure with properties of the emerging dynamics. Unlike traditional network analysis, which emphasizes static features such as node degree or clustering, our approach captures how behavior emerges over time through interactions across the network.
@@ 1.1 Complex quadratic dynamics
For iterations of a single quadratic map, one of the most striking results is that the trajectory (or orbit) of the critical point $z=0$ under iterations of the map captures global information on the behavior of all other orbits. The property of the critical orbit being bounded is particularly important in the context of long-term iterations. The Mandelbrot set appeared in the 1960s from the need to obtain an overarching description of this property over the whole family of quadratic maps. It is defined as the set of all $c\in\mathbb{C}$ for which the critical orbit of the complex quadratic function $f_{c}(z)=z^{2}+c$ is bounded. This is a fractal with fascinating characteristics that have been extensively analyzed since its introduction. The topological organization of the Mandelbrot set into hyperbolic bulbs (leading to its shape and fractal geometry) classifies all possible combinatoric behaviors that orbits can have in the family $f_{c}$. This makes the Mandelbrot set both a staple in discrete dynamical systems and a canonical representation for more general phenomena.
@@ 1.2 Complex quadratic maps on networks
Our work extends this framework from its original scope of single map iterations to the broader context of complex dynamical networks (CGNs). More specifically, we consider networks in which the coupled nodes are identical complex quadratic maps (CQNs). Quadratic maps were chosen for their historical importance and relative simplicity, both of which are desired properties in this proof-of-principle demonstration of the concept. Alternative families, such as higher-degree polynomials or rational maps, introduce additional critical points and poles, which make the associated theory substantially more complicated. Future studies will extend this approach to a broader class of discrete models for network nodes, with the aim of identifying both differences and universal properties across families.
The overall evolution of this discrete system takes the form of an iteration in $\mathbb{C}^{n}$, which acts in each component as
@@
The coupling is specified by the binary adjacency matrix $A\in\mathbb{R}^{n\times n}$, whose binary entries $A_{ij}$ mark an edge between nodes $i$ and $j$ in the graph, and signed weights $g\in\mathbb{R}^{n\times n}$, whose entries $g_{ij}$ provide the weight of the corresponding edge. Since nodes have the same intrinsic dynamics dictated by the parameter $c$, the emerging network dynamics are a direct reflection of the value of $c$ and of the specific $(A,g)$ network architecture at play. By analogy to the single map case, we define the equi-M set of the network as the set of all $c\in\mathbb{C}$ for which the critical point $0\in\mathbb{C}^{n}$ remains bounded under iteration. In this generalized context, many of the results on the properties and structure of the traditional Mandelbrot set are lost. Our theoretical work shows that the equi-M set no longer provides a full combinatorial description of the asymptotic behavior of the map [12]. However, our results strongly suggest that equi-M sets have quantifiable topological and fractal properties that reflect the network architecture, making them good candidates for testing and classifying how dynamic patterns evolve for different architectural profiles. In other words, the equi-M set still captures enough of the long-term dynamic behavior to be used as an assessment of the network’s dynamic range.
@@ 2.1 Dataset description
Our analysis is based on tractography-derived and functional neural connectivity data obtained from the S1200 Q4 release of the Human Connectome Project (HCP) [20]. Among others, the data contains Magnetic Resonance Imaging (MRI) scans from approximately 1200 healthy young adults between the ages of 20-40 balanced for sex, who underwent extensive resting-state, task, structural, and diffusion MRI. Conducting the current preliminary analysis on the whole dataset would have posed significant computational challenges. Therefore, this study analyzed a subgroup of 48 subjects, previously used by one of the authors in a statistical network study [21, 22]. The subjects were chosen to belong to the same batch for both functional and diffusion MRI, ensuring consistency in scanner protocols or software version that might affect the imaging. Hence, the subjects are a representative random subsample of the HCP dataset, comparable in size to our previous statistical studies on Structural connectomes. This choice is optimal for a proof-of-principle study, as it ensures computational feasibility without compromising statistical power. Furthermore, it can be expanded to the larger data set, if needed for validation and testing replicability of the study.
@@ 2.2 Network measures
(BCBC) is a global network measure of how often a node lies on the shortest path between all pairs of nodes in the network.
@@
(ASAS) measures how likely nodes in a network are to connect with other nodes that have the same degree.
@@
(GEGE) is the average (over all node pairs in the network) of the shortest path length between those two nodes.
@@
(TRTR) is the ratio of “triangles to triplets” in the network [30].
@@
($M$) are patterns of network connections among a specified number of nodes forming a sub-graph $g$, which act as building blocks for complex networks [30]. Given that our Structural connectivity matrices are undirected, our analysis focused on two types of three-node motifs: 113322 ($M1$, two links) and 113322 ($M2$, triangle).
@@
Graph-theoretical measures for Functional connectomes. A common approach to analyzing signed networks, particularly in the context of brain connectomes, is to separate the positive and negative sub-networks and examine them independently. More specifically, the two sub-networks are constructed using the modified weights
$g_{kj}^{+}=\begin{cases}g_{kj},&\quad g_{kj}\geq 0,\\ 0,&\quad\text{otherwise},\end{cases}\quad\text{and}\quad g_{kj}^{-}=\begin{cases}g_{kj},&\quad g_{kj}<0,\\ 0,&\quad\text{otherwise}.\end{cases}$
This method assumes that nodes exhibiting strong positive correlations (highly co-activated regions) and those with negative correlations (anti-correlated activity) form distinct functional communities. By analyzing these sub-networks separately, investigators aim to capture the intrinsic dynamics of cooperative and opposing activity patterns within the network [32]. However, this approach implicitly assumes that positive and negative sub-networks can independently encode key aspects of the network’s emergent behavior, an approach that has faced scrutiny [33, 34]. In this study, we aim to examine the validity of this assumption by exploring the extent to which overall network dynamics are influenced not only by these segregated sub-networks, but also by the intricate interplay between them. Therefore, for Functional networks, we separately compute the measures for the positive / negative sub-networks and also for the absolute value of the full functional connectome. We will call these the networks associated with the connectome.
@@ 2.3 Construction of equi-M sets
The equi-M sets are computed numerically for the connectome of each of the 48 subjects using the netbrot visualizer [35]. Each group of equi-M sets is rendered in an appropriately scaled bounding box with a resolution of $1200\times 1200$ pixels. The iteration proceeds for a maximum of $512$ steps using an escape radius $R_{e}=100$ until it decides if the chosen point is in the set. The chosen parameters give reasonably accurate numerical representations for the theoretical equi-M set. Illustrations of the corresponding equi-M sets for all 48 subjects are included in Figures˜12, 13 and 15, for all Structural and Functional connectomes that were used in this analysis.
@@
For each of the three types of connectomes (Structural, Rest, and Emotion) we computed statistical equi-M sets over all subjects in the data set. For each parameter point $c$ in the corresponding bounding box, we computed the fraction $f$ of the number of connectomes in the group, such that $c$ is in the equi-M set for that connectome. We illustrate these “stochastic” connectomes using pcolor plots, in which the “hot” color gradient (from yellow to black) represents the fraction $f$. We also took a mean-field approach to the data, and we generated equi-M sets for prototypical connectomes, computed as the mean of all connectomes in a data set (group average). These were defined in previous work [12] and are described in more detail in Section˜3.
@@ 2.4 Geometric measures of equi-M sets
The equi-M sets for all Structural and Functional connectomes in our dataset are shown in Figures˜12, 13 and 15 in order to allow a general view of their diversity in size and geometry and to facilitate direct visual comparisons. We use a variety of approaches to capture and quantify these visual differences and to understand their relationship to the network architecture.
@@
Figure 1: Measurable differences in equi-M set shape can be captured by computing a set of geometric measures, illustrated here for two random Structural connectomes in our data set: subject #22 (panel A) and subject #5 (panel B). In the left panel, we have the following landmarks and distances: $C=0.13$, T=0.9T=-0.9, $L=0.22+0.16i$, R=0.06+0.43iR=-0.06+0.43i, $d_{TC}=1.03$, $d_{TR}=0.94$, $d_{TL}=1.14$, $d_{CL}=0.19$, $d_{R\bar{R}}=0.41$, $\epsilon=0.87$. In the right panel: $C=0.08$, T=0.39T=-0.39, $L=0.13+0.06i$, R=0.03+0.2iR=-0.03+0.2i, $d_{TC}=0.47$, $d_{TR}=0.42$, $d_{TL}=0.53$, $d_{CL}=0.008$, $d_{R\bar{R}}=0.87$, $\epsilon=0.84$.
@@
Figure 2: Measurable differences in equi-M set shape captured by computing geometric measures for sets derived from Functional connectomes. The two panels illustrate the different geometry landmarks and corresponding measures between two Emotion connectomes: subject #28 (panel A) and subject #3 (panel B). In the left panel, we have the following landmarks and distances: C=0.01C=-0.01, $T=0.024$, $S=0.025+0.001i$, L=0.021+0.003iL=-0.021+0.003i, $R=0.003+0.026i$, $d_{TC}=0.04$, $d_{SR}=0.03$, $d_{SL}=0.04$, $d_{CL}=0.006$, $d_{R\bar{R}}=0.05$, $\epsilon=1.3$. In the right panel: C=0.008C=-0.008, $T=0.015$, $S=0.015$, L=0.013+0.007iL=-0.013+0.007i, R=0.0006+0.013iR=-0.0006+0.013i, $d_{TC}=0.024$, $d_{SR}=0.021$, $d_{SL}=0.03$, $d_{CL}=0.008$, $d_{R\bar{R}}=0.02$, $\epsilon=1.1$.
@@
For each individual equi-M set, we computed a collection of measures that capture details of the 2-dimensional position, size, and shape of the set. To begin with, we noticed that the Structural sets (corresponding to symmetric connectomes with all positive entries) replicate some broad topological features of the traditional equi-M set. This is not surprising, and is explained in more detail in [12]. In particular, the sets are symmetric about the real axis and exhibit a main cardioid that delimit the locus of $c$ for which the critical point converges to an attracting fixed point. A feature that appears to be robust among all Structural connectomes is that the main cardioid has a cusp on the right, with varying position and depth. The main cardioid extends to the left with a “tail” formation. However, the tail structure is not robust and presents high variability of shapes. The one consistent feature remains that the tip of the tail is the leftmost point of the set. Therefore, we used the position $C$ of the cusp (with $\operatorname{Im}(C)=0$) and the position $T$ of the tip of the tail (with $\operatorname{Im}(T)=0$) as landmarks for the stretch of the set along the real axis. In addition, we considered the coordinates of the point $R$ farthest from the real axis (with $\operatorname{Im}(R)>0$) and the coordinates of the rightmost points (the lip of the cusp) $L$ (with $\operatorname{Im}(L)>0$). To encode the general shape, we obtained the horizontal diameter $d_{TC}=|T-C|$, the vertical diameter $d_{R\bar{R}}=2|\operatorname{Re}(R)|$, the depth of the cusp as $d_{CL}=|C-L|$, the slant diameters from the tail to the highest point $d_{TR}=|T-R|$ and to the lip of the cusp $d_{TL}=|T-L|$, and the “elongation” of the set, as the ratio $\epsilon=d_{R\bar{R}}/d_{TC}$. The collection of these measures is expected to encompass key information about variability in shape among Structural equi-M sets (as suggested by Figure˜1, in which the measures are illustrated for two example sets).
@@
To identify similar landmarks for the Rest connectomes, we had to take into consideration the impact of the negative signs in the Functional connections on the topological structure. The first striking difference was that the orientation of the sets was “reversed” from that of the Structural connectomes. While all Rest sets still had a main cardioid body and a visually identifiable cusp, the cusp was left facing, and the (less prominent) tail formation was extending to the right. These differences are encouraging, as they indicate from the outset that the dynamics are qualitatively distinct in positive compared to signed networks. As before, we identified the cusp $C$, the tip of the tail $T$, the highest point $R$ and the upper lip of the cusp $L$ for Rest connectomes. In addition, we identify a separate marker $S$, which is the rightmost point in the set, similarly to $L$ in the Structural equi-M sets. Together with this new marker, we define a slant diameter $d_{SL}=|S-L|)$ and the depth of the tail $d_{ST}=|S-T|$. We used this expanded set of measures to characterize the shape of the Rest sets.
@@
The Emotion sets presented with even more variability in the shape of the main body (see the complete illustration of all sets in Figure˜15). Although most cardioids still had a left cusp, this was replaced in some cases by a more complex formation. Similarly, while most sets had a right tail, this was “stumpy,” and in some cases missing or replaced by another cusp. While in our theoretical work we are trying to analytically understand the conditions behind creating main cardioid cusps, this is beyond the scope of this paper. In this project, we simply use the same set of landmarks $(C,T,R,L,S)$ as for the Rest connectomes, and the same combinations between them, to broadly estimate the shape of the sets (as suggested by Figure˜2, in which the measures are illustrated for two example sets). However, in our Results and Discussion, we further comment on the significance of the differences in cusp and tail orientation and structure between the three types of connectomes.
@@
Our simulations suggest dramatic between-subject differences in the detail and variability of the equi-M set boundary, beyond the broad geometric measures described in the previous paragraphs. To capture these subtler properties, we constructed a truncated Fourier representation of the equi-M set boundary. For each equi-M set, the boundary was first identified and approximated using the Ramer–Douglas–Peucker algorithm [36] (part of the OpenCV library) to a specified tolerance of $10^{-4}P$, where $P$ represents an approximation of the set perimeter. The resulting polygonal boundary approximation was then used to obtain a truncated Fourier representation (as seen in Figure˜3). Note that the boundary of an equi-M set is fractal in nature and its Fourier representation does not always converge in the limit. Even so, investigating the resulting Fourier modes (i.e. coefficients in the discrete Fourier expansion) can provide valuable geometric information about the sets. Intuitively, lower-order modes capture the broad shape and orientation of the set, while higher-order modes capture the fine-grained complexity and fractality of the boundary.
We used the Fourier representation (see [37, Chapter 6]) of the boundary of the equi-M set to compute several additional measures. In general, the measures are invariant under translations, rotations, and (sometimes) scaling. Among these are the centroid $G$, the area $A$, and the perimeter $P$. Other measures, such as (average) curvature, are not representative due to the presence of cusps. Normalized Fourier modes (also known as Fourier descriptors) can also offer further insight into the large scale structure of the equi-M sets [37, Chapter 11].
@@
Figure 3: Fourier mode reconstruction of the Structural connectome from Figure˜1a (top) and Functional connectome from Figure˜2b (bottom) with $N+1$ modes in $\{-N/2,\dots,N/2\}$.